Complete the following items (carry out your answers to two decimal places):

As the junior public health epidemiologist in a large state in the Midwest Rust Belt, you have been asked by your supervisor to assist her in updating the statistics on opioid deaths. She is also asking for your help in writing a brief report describing the current situation for the annual report on drug abuse.

  1. In 2015, there were 3050 drug-related deaths, of which 2950 involved opioids. What percent of the deaths were due to opioids? Express your answer in the following format: XX.XX% (e.g., 54.81%).

  1. Calculate the ratio of opioid : non-opioid deaths. Express your answer as a ratio relative to 1 (e.g., 40.96 : 1).

  1. In 2010, the number of drug-related deaths was 1544. What is the percentage increase in drug deaths between 2010 and 2015? Express your answer in the following format: XX.XX% (e.g., 54.81%).

  1. How many deaths occurred per day on average in 2010 and in 2015?   
  1. Because the population size may change over time, you also decide to look at rates, which for drug overdose are generally expressed as deaths/100,000 residents. The state population for 2015 was estimated at 11,620,000. Calculate the rate of drug overdose deaths.

  1. The rate for 2010 was 13.70/100,000. What is the ratio between the 2015 and 2010 rates? Express your answer as a ratio relative to 1 (e.g., 40.96 : 1), with 2015 first.

  1. Express this change in words:
Answer #1

In 2015, there were 3050 drug-related deaths, of which 2950 involved opioids. What percent of the deaths were due to opioids? Express your answer in the following format: XX.XX% (e.g., 54.81%).

In 2015, Deaths due to opioids = 2950

       In percentage = 2950 / 3050 X 100 = 96.72%

Calculate the ratio of opioid : non-opioid deaths. Express your answer as a ratio relative to 1 (e.g., 40.96 : 1).

Deaths due to opioids = 2950

Deaths due to non opioids = 3050-2950 = 100

Ratio of opioid : non-opioid deaths = 2950:100 or 29.50:1

In 2010, the number of drug-related deaths was 1544. What is the percentage increase in drug deaths between 2010 and 2015? Express your answer in the following format: XX.XX% (e.g., 54.81%).

In 2010, number of drug-related deaths = 1544

In 2015, number of drug-related deaths = 3050

Difference in deaths number between 2010 and 2015= 3050-1544 =1506

Total number of deaths in 2010 and 2015 = 3050+1544 = 4594

Percentage increase in drug deaths between 2010 and 2015 = 1506/4594 X 100 =32.78%

How many deaths occurred per day on average in 2010 and in 2015?   

In 2010, number of drug-related deaths = 1544

Deaths occurred per day on average in 2010 = 1544/365days = 4.23 deaths per day

In 2015, number of drug-related deaths = 3050

Death occurred per day in 2015 = 3050/365 = 8.35 deaths per day

Because the population size may change over time, you also decide to look at rates, which for drug overdose are generally expressed as deaths/100,000 residents. The state population for 2015 was estimated at 11,620,000. Calculate the rate of drug overdose deaths.

State population for 2015 = 11,620,000

Number of drug-related deaths in 2015, = 3050

Death rate is described as number of deaths recorded is divided by the number of people in the population, and then multiplied by 100000 (as given in question)

Rate of drug overdose = 3050/11620000 X 100000 =26.25 per 100,000 residents

The rate for 2010 was 13.70/100,000. What is the ratio between the 2015 and 2010 rates? Express your answer as a ratio relative to 1 (e.g., 40.96 : 1), with 2015 first.

Rate of drug overdose for 2015 = 26.25/100,000

Rate of drug overdose for 2010 = 13.70/100,000

Ratio between the 2015 and 2010 rates = 26.25 : 13.70 (by division) = 1.92 : 1

Express this change in words:

In 2015, Deaths due to opioids were 96.72% and ratio of opioid : non-opioid deaths was 29.50:1. Drug related deaths between 2010 and 2015 were increased by 32.78%. The average number of Death occurred per day between 2010 and 2015 was also almost doubled (increased) from 4.23 deaths per day to 8.35 deaths per day. A rate of 26.25 drug overdose per 100,000 residents was observed in 2015 with a ratio of 1.92 :1 in comparison to 2010. Overall, death number increased in 2010 due to increased consumption in population .